OpenStudy (anonymous):

Given f(x) = x + 1 and g(x) = x – 1, find (fg)(x). is it x^2-1

OpenStudy (solomonzelman):

\(\large\color{black}{ f(x) = x + 1 }\) \(\large\color{black}{ \color{green}{ g(x)} = \color{red}{ x- 1} }\) \(\large\color{black}{ f(\color{green}{ g(x)} ) = (\color{red}{ x- 1} ) + 1 }\) \(\large\color{black}{ f(\color{green}{ g(x)} ) = ? }\)

jimthompson5910 (jim_thompson5910):

SolomonZelman I think it might be "f times g"

OpenStudy (anonymous):

@jim_thompson5910 is right its f times g

jimthompson5910 (jim_thompson5910):

so you'd do (f*g) = f(x)*g(x) = (x+1)*(x-1) = x^2 - 1 which means you are indeed correct

OpenStudy (anonymous):

Thank you (:

jimthompson5910 (jim_thompson5910):

you're welcome